Table 1
22.8 27.0 27.9 30.4 33.4
24.8 27.2 27.9 31.1 33.9
24.8 27.4 28.2 31.4 35.3
26.0 27.4 29.4 32.4 35.7
26.0 27.4 29.6 33.1 36.3
26.1 27.6 29.8 33.2 40.2

a) Obtain the suitable number of classes and the class width of the given data.

b) Develop a frequency distribution table for the data in Table 1. Take 22.8 as the lower limit of the first class.
1
Expert's answer
2019-05-13T12:56:52-0400
"22.8,\ 24.8,\ 24.8,\ 26.0,\ 26.0,\ 26.2,""27.0,\ 27.2,\ 27.4,\ 27.4,\ 27.4,\ 27.6,""27.9,\ 27.9,\ 28.2,\ 29.4,\ 29.6,\ 29.8,""30.4,\ 31.1,\ 31.4,\ 32.4,\ 33.1,\ 33.2,""33.4,\ 33.9,\ 35.3,\ 35.7,\ 36.3,\ 40.2"

"n=30, number\ of\ classes=\sqrt{n}"

a) We can use 6 classes.


b) Develop a frequency distribution table for the data in Table 1. Take 22.8 as the lower limit of the first class.

"\begin{matrix}
 22.8\text{---}25.7 & \ 3 \\
 25.8\text{---}28.7 & 12
\end{matrix}""\begin{matrix}
 28.8\text{---}31.7 & \ 6 \\
 31.8\text{---}34.7 & \ 5
\end{matrix}""\begin{matrix}
 34.8\text{---}37.7 & \ 3 \\
 37.8\text{---}40.7 & \ 1
\end{matrix}"